# Demonstration of quadrature interferometric metrology of translation and tilt: QUIMETT

### v1

@@ 4.1 Tilt Measurement
We hereafter call those interference conditions, mid(--), dark, mid($+$), and bright. The interference condition was adjusted by the piezo actuator on the reference mirror.
@@
During the measurements, the tilt of the test mass simulator was excited by triangular waves at 1 Hz with an amplitude of 10 $\mu$rad. The triangular waves were generated by a function generator and sent to the set of piezo actuators driving the test mass simulator as shown in Fig. 2. The pitch and yaw excitations are applied one at a time.
The results of tilt measurements for yaw and pitch directions in the bright fringe are shown in Fig. 3 together with the theoretical values for comparison. The measured voltage data are calibrated to the tilt data using Eq. (14). In this calculation, we assume $R_{\mathrm{PBS2}}=1$ and $T_{\mathrm{PBS2}}=1$ and we use the parameters shown in Tab. 2. In both yaw and pitch directions, the measured values are found to be almost half of the theoretical expectation values. However, the tilt motions are measured consistently during the measurement time. Some discrepancy could be due to the uncharacterized optical losses in the optics in the actual setup. The tilt measurement and residual show the same waveform, indicating that there is not large distortion in the measurement but a difference in scaling factor.
@@
Gouy phase ($\eta_{\mathrm{ITF}}+\eta_{\mathrm{QPD}}$) 106 deg
@@
Beam size on the actuated mirror 116 $\mu$m
@@ 4.2 Tilt measurement noise
Figure 5: Tilt noises in amplitude spectral densities. Measurements are performed at four interference conditions. (b) shows bright fringe, (c) shows mid($+$), (d) shows dark fringe, and (e) shows mid(--). The measurement results without the piezoelectric element (a) are also shown to evaluate the effect of the piezoelectric element.
In addition, an extra measurement was performed with the mirror in the optical system replaced with a mirror mounted on a standard mount to see the influence from the piezo-actuated mirror. The data is shown with label (a) in Fig. 5.
In any interference conditions, sharp features are seen at frequencies above 10 Hz. These noise features increases in the mid fringes. On the other hand, we find the noise levels to be almost identical between the different interference conditions at frequencies below 10 Hz. We determined this noise to be the digitization noise from the data recorder as shown in Fig. 6. In fact, the tilt sensing noise above 10 Hz is almost entirely dominated by the digitization noise.
@@ 5 Discussions
Besides the demonstration of the tilt measurement, there are several critical considerations which have to be addressed before adopting QUIMETT as the local sensor for the drag-free control. First, a further look into the length-to-tilt coupling is necessary. As described in Sec. 4, the response to tilts did not change as the function of the interference condition. However, there could be unexpected cross-coupling where the dynamical motion in the length degree of freedom leak into the tilt signals. We were not able to quantify such couplings in this setup because the piezo-actuated mirror showed an inherent cross-coupling between the tilts and length presumably due to the mechanical imbalance.
Similarly, the tilt-to-length coupling must be studied. As shown in Eq. (9), the length signal is immune to the variation in the light intensity. This holds true as long as the responsibities of the PD and QPDs do not vary. However, the total amount of light detected by the QPDs could easily be a function of the beam spot displacement caused by the test mass tilts because the QPDs are inherently more sensitive to this coupling due to the fact that there are gaps between the segments where the light is not converted into the photocurrent. Such a tilt-to-length coupling can also occur in the single-element PD due to the inhomogeneous responsibity across the PD surface area, but it is typically smaller than that of the QPDs.
Furthermore, comparing against the single-element PDs, the use of QPDs tends to reduce the signal-to-noise ratio for length sensing due to the presence of the gap between the segments. This could lead to a situation where the level of noise floor in the length measurement varies depending on the spot position on QPDs. This may impact on the background noise estimation in gravitational wave data analysis where the stationary noise over a long time scale is often preferable.
@@
As shown in Eqs. (10) and (11), each QPD signal contains not only the tilt component but also the one associated with the length difference or $l_{x}-l_{y}$. Consequently the electronics and measurement instrument have to deal with a wide range of signals. In the current setup, the tilt signals of each QPD were recorded by the measurement instrument and subsequently the sum of the two QPD signals was computed in the post process. This made the contribution of ADC noise particularly large due to the relatively narrow dynamic range in the ADC. Therefore, we can improve the noise floor by introducing analog electronics such that the summing process is completed before the ADC where the ADC range can be adjusted to be a smaller value.
@@
Finally, refinement of the optical design to reduce the volume and weight for the space application is necessary. Equation (14) reveals that optical path length does not influence the tilt sensitivity. This means the reduction of the optical path lengths enables a design with smaller volume without compromising the sensitivity. In particular, shortening the distance between PBS1 and NPBS in the Fig. 1 is considered to be effective for achieving a compact design.
@@ 6 Conclusion
This paper proposes a new local sensor concept for the drag-free control in space. The sensor is based on the quadrature interferometer and named QUIMETT. QUIMETT enables the simultaneous readout of both translation and tilt degrees of freedom using a single interfermeter module. This paper presented for the first time the working principle of the tilt measurements and its demonstration by building the dedicated optical system. In the experimental verification, the main focus was put on the demonstration of the tilt readout which were designed to be independent of the interference condition. The experimental demonstration successfully confirmed this function. The tilt sensitivity currently marks $10$ nrad/Hz1/2 at $0.1$ Hz and did not change in four different interference conditions as expected. We discussed the studies which have to be further looked in before considering this scheme as a viable solution for the drag-free control.
@@ Acknowledgment
We thank Advanced Machining Technology Group and K. Komori for technical support, thank C. M. Mow-Lowry and S. Sato for useful discussions. This project was supported by JSPS KAKENHI Grant Number JP20J01928 and JP22K14067. K.N. is an employee of LQUOM, Inc.
@@
K.N. and K.M. carried out the experiment. K.N. conceived the original idea. K.M. and K.I. developed the theoretical formalism, and performed the analytic calculations. K.I. supervised the project. All authors discussed the results and contributed to the final manuscript. K.N. and K.M. contribute equally the presented work.
@@ Appendix A Mathematical foundation
Here, we show the mathematical background of the optical analysis in this paper. Upon transmission and reflection, the electric field in a polarization acquire sets of coefficients given matrix form as,
@@
where diag represents the diagonal matrix, $T$ and $R$ are the transmittance and reflectance of the optic, respectively. Using these matrices, one can express the polarizing beam splitter (PBS) in the $6\times 6$ format as
@@
We deal with misalignment of optical element [25] as an additive matrix as
@@
where $\Theta_{A}$ represents the normalized tilt defined by $\Theta_{A}=(\theta_{A}\pi\omega(z))/\lambda$ with $\theta_{A}$ a rotation in the optic, $\lambda$ the wavelength and $w(z)$ the beam radius at the propagation point $z$. Eq. (19) is only valid when $\Theta_{\textrm{A}}$ is sufficiently small. The propagation for by optical path transmission is represented as
@@
The half wave plate can be expressed as [26]
@@
with $\phi$ the angle between the fast axis to half wave plate. Jones matrix of quarter wave plate is:
@@
where $\psi$ is angular of quarter wave plate to fast axis.
@@
For tilt sensing by the QPDs, using the laser intensity in the hemisphere of the QPD. For detect tilt in the yaw direction, the laser intensity in the left and right half is discussed here. We hereafter consider one polarization mode since, in the setup of QUIMETT, the QPDs detect one polarized light due to the PBS in front of the QPDs. Laser intensity of the right half of the QPD is denoted as
@@
where $x$ represents horizontal position on the QPD, and $U_{00}$ and $U_{10}$ ($U_{01}$) represent the fundamental and first-order Hermite polynomial denoting the horizontally (vertically) spatial mode of the Gaussian-Hermite beam, respectively. Here, we assume that $|E_{00}|\gg|E_{10/01}|$ and the size of the QPD is much larger than the beam size on the QPD. $U_{00}$ and $U_{10}$ are defined as
@@
where $\omega$ and $R$ are beam size and curvature of the electric field on the QPD, respectively. From this definition, we can obtain the following equation,
@@
Tilt sensing by QPD is provided by subtraction of one hemisphere from the other,
@@ Appendix B Effect of reference mass
$\begin{split}\mathbb{E}^{\textrm{QPD}n}=&\mathbb{F}(n)\mathbb{W}^{\left(4\right)}\left(\frac{\pi}{4}\right)\mathbb{P}(l_{3},\eta_{3})\mathbb{A}_{\textrm{t}}\left(\frac{1}{2}\right)\mathbb{P}(l_{1},\eta_{1})\mathbb{W}^{\left(2\right)}\left(-\frac{\pi}{8}\right)\\ &\times\{\mathbb{B}_{v}\left(R_{\textrm{PBS2}}\right)\mathbb{P}(l_{y},\eta_{\mathrm{ITF}})\mathbb{M}\left(\Theta_{\textrm{yaw}},\Theta_{\textrm{pitch}}\right)\mathbb{P}(l_{y},\eta_{\mathrm{ITF_{in}}})\mathbb{B}_{v}\left(R_{\textrm{PBS2}}\right)\\ &+\mathbb{B}_{h}(T_{\textrm{PBS2}})\mathbb{P}(l_{x},\eta_{\mathrm{ITF}})\mathbb{M}\left(\Theta_{\textrm{yaw2}},\Theta_{\textrm{pitch2}}\right)\mathbb{P}(l_{x},\eta_{\mathrm{ITF_{in}}})\mathbb{B}_{h}\left(T_{\textrm{PBS2}}\right)\}\\ &\times\mathbb{W}^{\left(2\right)}\left(\frac{\pi}{8}\right)\mathbb{P}(l_{1},\eta_{1_{\mathrm{in}}})\mathbb{A}_{\textrm{r}}\left(\frac{1}{2}\right)\mathbb{E}_{\textrm{in}}.\end{split}$ (33)

### v2

@@ 4.1 Tilt Measurement
We hereafter call these interference conditions, mid(--), dark, mid($+$), and bright. The interference condition was adjusted by the piezo actuator on the reference mirror.
@@
During the measurements, the tilt of the test mass simulator was excited by triangular waves at 1 Hz with an amplitude of 10 $\mu$rad. The triangular waves were generated by a function generator and sent to the set of piezo actuators driving the test mass simulator as shown in Fig. 2. The pitch and yaw excitations were applied one at a time.
The results of tilt measurements for yaw and pitch directions in the bright fringe are shown in Fig. 3 together with the theoretical values for comparison. The measured voltage data are calibrated to the tilt data using Eq. (14). In this calculation, we assume $R_{\mathrm{PBS2}}=1$ and $T_{\mathrm{PBS2}}=1$ and we use the parameters shown in Tab. 2. In both yaw and pitch directions, the measured values are found to be almost half of the theoretical expectation values. However, the tilt motions are measured consistently during the measurement time. Some discrepancy could be due to the uncharacterized optical losses in the optics in the actual setup. Further discussion is provided in Section 5. The tilt measurement and residual show the same waveform, indicating that there is not large distortion in the measurement but a difference in the scaling factor.
@@
Gouy phase ($\eta_{\mathrm{ITF}}+\eta_{\mathrm{QPD}}$) $124$ deg
@@
Beam diameter on the actuated mirror 340 $\mu$m
Beam diameter on the QPDs 700 $\mu$m
@@ 4.2 Tilt measurement noise
Figure 5: Tilt noises in amplitude spectral densities. Measurements were performed at four interference conditions. (b) shows bright fringe, (c) shows mid($+$), (d) shows dark fringe, and (e) shows mid(--). The measurement results without the piezoelectric element (a) are also shown to evaluate the effect of the piezoelectric element.
In addition, an extra measurement was performed with the mirror in the optical system replaced with a mirror mounted on a standard mount to see the influence from the piezo-actuated mirror. The data are shown with label (a) in Fig. 5.
In any interference conditions, sharp features are seen at frequencies above 10 Hz. These noise features increase in the mid fringes. On the other hand, we find the noise levels to be almost identical among the different interference conditions at frequencies below 10 Hz. We determined this noise to be the digitization noise from the data recorder as shown in Fig. 6. In fact, the tilt sensing noise above 10 Hz is almost entirely dominated by the digitization noise.
@@ 5 Discussions
Besides the demonstration of the tilt measurement, there are several critical considerations which have to be addressed before adopting QUIMETT as the local sensor for the drag-free control.
5.1 Comparison between experimental and theoretical results
Figures 3 and 4 show that the theoretical predictions and experimental measurements exhibit a discrepancy of approximately a factor of two. The tilt angle of the movable mirror has been independently verified using an optical lever within 1% error. Therefore, we discuss systematic errors in other parts of the system as the possible source of the discrepancy.
Our analysis suggests that this discrepancy is not caused by a single major error. Instead, it results from the accumulation of several small systematic errors. These include minor deviations in mirror reflectivity, absorption losses in waveplates, imperfections in the extinction ratio of the PBS, variations in the quantum efficiency of the QPDs, and discrepancies in the transimpedance amplifier gain. Although each of these factors may have a limited individual impact, their combined effects contribute to the observed mismatch. Reducing these uncertainties is crucial for improving the accuracy of the system. Addressing these issues will be an important objective for future work. Nevertheless, these systematic errors influence all fringe conditions in the same manner. As a result, the consistency of our experimental results across different fringe conditions suggests that their effects are uniform. This indicates that the proposed method successfully integrates tilt-sensing functionality into a quadrature interferometer. At the same time, it maintains the capability for displacement sensing over a wide range.
5.2 Noise performance evaluation
We also evaluated the noise performance of QUIMETT. Compared to conventional optical lever tilt sensors [26], our current setup exhibited a noise level approximately ten times higher. The primary reason for this increased noise is the lack of full optical optimization, particularly in terms of beam waist positioning and the collimator and lens configuration.
@@ 5.2 Noise performance evaluation
A major source of noise in the experiment was digitization noise from the ADC since, as shown in Eqs. (10) and (11), each QPD signal contains not only the tilt component but also the one associated with the length difference or $l_{x}-l_{y}$. Consequently the electronics and measurement instrument have to deal with a wide range of signals. In the current setup, the tilt signals of each QPD were recorded by the measurement instrument and subsequently the sum of the two QPD signals was computed in the post process. This made the contribution of ADC noise particularly large due to the relatively narrow dynamic range in the ADC. Therefore, we can improve the noise floor by introducing analog electronics such that the summing process is completed before the ADC where the ADC range can be adjusted to be a smaller value.
5.3 Coupling effects between translation and tilt
A further look into the length-to-tilt coupling is necessary. As described in Sec. 4, the response to tilts did not change as a function of the interference condition. However, there could be unexpected cross-coupling where the dynamical motion in the length degree of freedom leaks into the tilt signals. We were not able to quantify such couplings in this setup because the piezo-actuated mirror showed an inherent cross-coupling between the tilts and length presumably due to the mechanical imbalance.
Similarly, the tilt-to-length coupling must be studied. As shown in Eq. (9), the length signal is not affected by variations in light intensity, as long as the efficiencies of the PD and QPDs remain constant. However, the total amount of light detected by the QPDs could easily be a function of the beam spot displacement caused by the test mass tilts because the QPDs are inherently more sensitive to this coupling due to the fact that there are gaps between the segments where the light is not converted into the photocurrent. Such a tilt-to-length coupling can also occur in the single-element PD due to the inhomogeneous responsivity across the PD surface area, but it is typically smaller than that of the QPDs.
Furthermore, comparing against the single-element PDs, the use of QPDs tends to reduce the signal-to-noise ratio for length sensing due to the presence of the gap between the segments. This could lead to a situation where the level of noise floor in the length measurement varies depending on the spot position on the QPDs. This may impact on the background noise estimation in gravitational wave data analysis where the stationary noise over a long time scale is often preferable.
5.4 Optimization and future prospects
For future applications of QUIMETT, several optimizations must be considered. Here, we consider the measurement range. The measurement range of our device is determined by the normalized tilt parameter $\Theta$ which was treated as a small second-order effect in Eqs. (7) and (8). This suggests that the current dynamic range is approximately $\sim 1$ mrad, where $\Theta^{2}$ remains on order of magnitude of unity. For future applications, such as the potential use of QUIMETT in DECIGO, we estimate the required measurement range by considering a test-mass module configuration similar to the LISA Pathfinder [27] extended to DECIGO [12]. Assuming a $4$ mm gap in the sensor module and a $30$ cm mirror size, the necessary range would be $\sim 10$ mrad. This exceeds the current capability, indicating that further design considerations, such as incorporating additional optical elements like lenses, will be necessary to extend the measurement range.
@@ 5.4 Optimization and future prospects
Finally, refinement of the optical design to reduce the volume and weight for the space application is necessary. Equation (14) reveals that optical path length does not influence the tilt sensitivity. This means the reduction of the optical path lengths enables a design with smaller volume without compromising the sensitivity. In particular, shortening the distance between PBS1 and NPBS in Fig. 1 is considered to be effective for achieving a compact design.
@@ 6 Conclusion
This paper proposes a new local sensor concept for the drag-free control in space. The sensor is based on the quadrature interferometer and named QUIMETT. QUIMETT enables the simultaneous readout of both translation and tilt degrees of freedom using a single interferometer module. This paper presents for the first time the working principle of the tilt measurements and its demonstration by building the dedicated optical system. In the experimental verification, the main focus was put on the demonstration of the tilt readout which was designed to be independent of the interference condition. The experimental demonstration successfully confirmed this function. The tilt sensitivity currently marks $10$ nrad/Hz1/2 at $0.1$ Hz and did not change in four different interference conditions as expected. We discussed the studies which have to be further looked in before considering this scheme as a viable solution for the drag-free control.
@@ Acknowledgment
We thank Advanced Machining Technology Group and K. Komori for their technical support. We also appreciate C. M. Mow-Lowry and S. Sato for their useful discussions. This project was supported by JSPS KAKENHI Grant Number JP20J01928 and JP22K14067. K.N. is an employee of LQUOM, Inc.
@@
K.N. and K.M. carried out the experiment. K.N. conceived the original idea. K.M. and K.I. developed the theoretical formalism, and performed the analytic calculations. K.I. supervised the project. All authors discussed the results and contributed to the final manuscript. K.N. and K.M. contributed equally to the present work.
@@ Appendix A Mathematical foundation
Here, we show the mathematical background of the optical analysis in this paper. Upon transmission and reflection, the electric field in a polarization acquires sets of coefficients given matrix form as,
@@
where diag represents the diagonal matrix, $T$ and $R$ are the transmittance and reflectance of the optical component, respectively. Using these matrices, one can express the polarizing beam splitter (PBS) in the $6\times 6$ form as,
@@
We deal with misalignment of an optical element [25] as an additive matrix as
@@
where $\Theta_{A}$ represents the normalized tilt. It is defined by $\Theta_{A}=(\theta_{A}\pi\omega(z))/\lambda$, where $\theta_{A}$ is a rotation in the optical element, $\lambda$ is the wavelength and $w(z)$ is the beam radius at the propagation point $z$. Equation (19) is only valid when $\Theta_{\textrm{A}}$ is sufficiently small. The propagation through an optical path transmission is represented as
@@
The half-wave plate is expressed as [28]
@@
where $\phi$ is the fast-axis angle of the half-wave plate. Jones matrix of quarter-wave plate is expressed as,
@@
where $\psi$ is the fast-axis angle of the quarter-wave plate.
@@
For tilt sensing by the QPDs, we consider the laser intensity in the hemisphere of the QPD. To detect tilt in the yaw direction, the laser intensity in the left and right halves is discussed here. We hereafter consider one polarization mode since, in the setup of QUIMETT, the QPDs detect one polarized light due to the PBS in front of the QPDs. The laser intensity of the right half of the QPD is denoted as
@@
where $x$ represents horizontal position on the QPD, and $U_{00}$ and $U_{10}$ ($U_{01}$) represent the fundamental and first-order Hermite polynomials denoting the horizontally (vertically) spatial mode of the Gaussian-Hermite beam, respectively. Here, we assume that $|E_{00}|\gg|E_{10/01}|$ and the size of the QPD is much larger than the beam size on the QPD. $U_{00}$ and $U_{10}$ are defined as
@@
where $\omega$ and $R$ are the beam size and the curvature of the electric field on the QPD, respectively. From this definition, we can obtain the following equation,
@@
Tilt sensing by the QPD is provided by subtraction of one hemisphere from the other,
@@ Appendix B Effect of reference mass
$\begin{split}\mathbb{E}^{\textrm{QPD}n}=&\mathbb{F}_{n}\mathbb{W}^{\left(4\right)}\left(\frac{\pi}{4}\right)\mathbb{P}(l_{3},\eta_{3})\mathbb{A}_{\textrm{t}}\left(\frac{1}{2}\right)\mathbb{P}(l_{1},\eta_{1})\mathbb{W}^{\left(2\right)}\left(-\frac{\pi}{8}\right)\\ &\times\{\mathbb{B}_{v}\left(R_{\textrm{PBS2}}\right)\mathbb{P}(l_{y},\eta_{\mathrm{ITF}})\mathbb{M}\left(\Theta_{\textrm{yaw}},\Theta_{\textrm{pitch}}\right)\mathbb{P}(l_{y},\eta_{\mathrm{ITF_{in}}})\mathbb{B}_{v}\left(R_{\textrm{PBS2}}\right)\\ &+\mathbb{B}_{h}(T_{\textrm{PBS2}})\mathbb{P}(l_{x},\eta_{\mathrm{ITF}})\mathbb{M}\left(\Theta_{\textrm{yaw2}},\Theta_{\textrm{pitch2}}\right)\mathbb{P}(l_{x},\eta_{\mathrm{ITF_{in}}})\mathbb{B}_{h}\left(T_{\textrm{PBS2}}\right)\}\\ &\times\mathbb{W}^{\left(2\right)}\left(\frac{\pi}{8}\right)\mathbb{P}(l_{1},\eta_{1_{\mathrm{in}}})\mathbb{A}_{\textrm{r}}\left(\frac{1}{2}\right)\mathbb{E}_{\textrm{in}}.\end{split}$ (33)